OpenStudy (anonymous):

solve the equation (7a/3a+3)-(5/4a-4)=(3a/2a+2)

OpenStudy (anonymous):

Is this the equation? \[\frac{ 7a }{ 3a+3 } - \frac{ 5 }{ 4a-4 } = \frac{ 3a }{ 2a+2 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a=3, a=-1/2

OpenStudy (anonymous):

can you show me how you got those answers please

OpenStudy (anonymous):

multiply 12 both side

OpenStudy (anonymous):

take the right term to left and second term on the left to right..

OpenStudy (anonymous):

\[\frac{ 28a }{ a+1 } - \frac{ 18a }{ a+1 }=\frac{ 15}{ a-1 }\]

OpenStudy (anonymous):

\[\frac{ 10a }{ a+1 } =\frac{ 15}{ a-1 }\]

OpenStudy (anonymous):

\[\frac{ 2a }{ a+1 } =\frac{ 3}{ a-1 }\]

OpenStudy (anonymous):

\[2a(a-1)=3(a+1) \]

OpenStudy (anonymous):

2a^2-2a=3a+3 2a^-5a-3=0 (2a+1)(a-3)=0 a=3 a=-1/2

OpenStudy (anonymous):

ok thank you i got it can you help me with this one 1=(1/1-a)+(a/a-1)

OpenStudy (anonymous):

i don't understand why you times everything by 12? I get that it's the least common denominator so are you trying to find the least common denominator of the numbers that have a variable after them or what?

OpenStudy (anonymous):

yes 12 is the lcm of 2,3,4. you don't need to do that, but this is one way to solve it and it makes our job easier, at least for me..